OpenStudy (anonymous):

Find the value of k so that the ordered pair satisfies the equation: 6x - ky = k; (2,2) A. 6 B. 9 C. 4 D. -4

OpenStudy (ajprincess):

Plug in x=2 and y=2 in the given equation and solve for k. Can u do @smartkidd18?

OpenStudy (anonymous):

Is it C?

OpenStudy (ajprincess):

yup:) 6*2-k*2=k 12-2k=k 12=3k 12/3=k k=4

OpenStudy (anonymous):

Thanks!

OpenStudy (ajprincess):

Welcome:)